OpenStudy (anonymous):

In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive. One point for the equation, one point for the solution.

OpenStudy (hoblos):

x=y+53 x+y=221 y+53+y=221 2y=168 y=84 x=84+53=137

OpenStudy (hoblos):

x= votes Connie received y= votes Carl received

OpenStudy (anonymous):

votes carl got = x then votes connie got = x + 53 x + x + 53 = 221 2x = 221 - 53 x = 168/2 x = 84 so carl got 84 votes connie got 84 + 53 = 137 votes

OpenStudy (anonymous):

thanks friends.....☺

OpenStudy (anonymous):

thanks!